OpenStudy (anonymous):

Four times a certain number increased by 6 is 94.

OpenStudy (anonymous):

94-6/4 i think

OpenStudy (triciaal):

let the unknown number = x times means multiply increased means add Four times a certain number increased by 6 is 94. 4x + 6 = 94

OpenStudy (anonymous):

either way you should get the correct answer

OpenStudy (triciaal):

solve for x if you want to find the number

OpenStudy (anonymous):

Four times a certain number increased by 6 is 94. 4(6x) = 94 4x + 6 = 94 4(x + 6) = 94 x + 6 = 4(94) B is wrong

OpenStudy (anonymous):

i think my website is wrong because b looks right

OpenStudy (anonymous):

the answer will be a decimal

OpenStudy (triciaal):

maybe you should use parenthesis the problem could be this 4 (x + 6) = 94

OpenStudy (anonymous):

thank yal

OpenStudy (triciaal):